OpenStudy (lilkg77):

Use the explicit formula an = a1 + (n - 1) • d to find the 350th term of the sequence below. 57, 66, 75, 84, 93, ... A. 3234 B. 3207 C. 3141 D. 3198

OpenStudy (danjs):

that type has terms that increase linearly, the same amount each time 'd'

OpenStudy (danjs):

you know the start term, and the difference to the next 'd', that is all you need

OpenStudy (michele_laino):

hint: the constant \(d\) of the series, is such that: \[d = 93 - 84 = 84 - 75 = 75 - 66 = 66 - 57 = ...?\]

OpenStudy (danjs):

an = a1 + (n - 1) • d n starts at 1, and results in a1, then every next term adds on an additional 'd' , any number term you want to know, it is just the first term value a1, and one less than the term number times the common difference 'd'

OpenStudy (anonymous):

a1 + (n-1) d /. { a1 -> 57, n -> 350, d -> 66-57 } /. means apply the replacement rules on the right to the expression on the left and then evaluate the result. What looks like mumbo jumbo on the right hand side is a list of replacement rules from the Mathematica computer program. /. is the Mathematica replacement operator. Did you get 3198 ?

OpenStudy (lilkg77):

Yes thanks @robtobey @DanJS